# Enhanced Generative Data Augmentation for Semantic Segmentation via Stronger Guidance

### v2

@@ 15                  $\mathcal{D}_{gen}\leftarrow\mathcal{D}_{gen}\cup\{\mathcal{I}_{gen}\}$
In the default setting, we choose an appropriate $n_{gen}$ such that $|D_{gen}|$ approx $|D_{origin}|$ where |||| . |||| is the number of images in the dataset. In Section 4.3.2, we further discuss changing the number of synthetic images generated.
@@ 4 RESULTS
Figure 4: Some results show the limitations of direct generation and the ability of our method to overcome them.
@@ 4.1.1 Dataset
In this section, we evaluate our method on the segmentation datasets VOC7 and VOC12 [Everingham et al., 2015]. PASCAL VOC 2007 has 422 images annotated for semantic segmentation, split into 209 training and 213 validation images. Meanwhile, VOC12 has training and validation sets, including 1.464 and 1.449 images. In addition to training on the entire VOC12 dataset (1.464 images), we also train our model using 1/2 (732 images), 1/4 (366 images), 1/8 (183 images), and 1/16 (92 images) partition protocols [Wang et al., 2022]. Evaluating our method on these few-sample datasets demonstrates its effectiveness in real-world scenarios.
@@ 4.1.2 Implementation details
We construct our framework on the deep learning PyTorch framework [Paszke et al., 2019] and T2I-Adapter [Mou et al., 2023] using Stable Diffusion XL 1.0 [Podell et al., 2024] with 30 time steps. We generate data using values for $\omega_{1}$ and $\omega_{2}$, as defined in Section 4.3. For semantic segmentation, we employ the DeepLabV3+ [Chen et al., 2018], PSPNet [Zhao et al., 2017], and Mask2Former [Cheng et al., 2022] with segmenters implemented in the MMSegmentation framework [MMSegmentation Contributors, 2020]. We utilize the SGD optimizer with standard settings in MMSegmentation. We train our models with an input image size of 512$\times$512 on the Pascal VOC datasets, including VOC7 and VOC12. The performance of the trained model on the applied data is assessed using the Mean Intersection over Union (mIoU) metric.
@@ 4.2.1 Quantitative results:
The results presented in Table 1 demonstrate that combining augmented data with the original dataset improves the performance of various segmentation models on the VOC7 and VOC12 datasets. All models, including DeepLabV3+, PSPNet, and Mask2Former, significantly improve when augmented data. DeepLabV3+ consistently performs the best across different dataset sizes. We note that as the amount of real-world data increases, the accuracy of the generated images and labels becomes crucial. Mismatched generated images in the synthetic data can lead to performance degradation; this is observed in Mask2Former when trained on the VOC12 dataset with 1464 images.
4.2.2 Qualitative results
In Figure 4, each row presents three images: the original image, the image generated by the original generative model, the description generated by BLIP2, and the image generated by the generative model combined with our method. The first two rows illustrate cases where the labeled classes in the image are either missing from the description or incorrectly described. This causes the model to either fail to generate those classes correctly (row 1) or produce images with blurred or unrecognizable objects (row 2). In the last row, the image is rather simple, and the prompt includes all labeled classes, but the generated image is still incorrect. With our pipeline, all three images have significantly improved with certain objects being fully generated and correctly positioned. This demonstrates that our method can effectively overcome the limitations of directly using the controllable imaging model.
4.3 Ablation Study
4.3.1 Effects of different methods:
Table 2: Impact of Class-Prompt Appending (1), Visual Prior Combination (2), Class balancing algorithm (3), and Post Filter [Fang et al., 2024] (4).
@@ 4.3.1 Effects of different methods:
$\checkmark$ $\checkmark$ $\checkmark$ 49.72
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 51.23
@@
49.72
The performance of the proposed methods is summarized in Table 2. When using the Baseline with generated prompts, the performance is only as low as 42.15 % mIoU, which is lower than when training on the original data. This demonstrates the essential nature of the proposed methods when generating synthetic data. Our method yields a 49.72 % mIoU, showing its effectiveness. Additionally, we experimented with combining the Post Filter with the Category-Calibrated CLIP Rank [Fang et al., 2024] for generating synthetic data. The result of this combination is higher, which shows that our method can combine filters from previous studies [Fang et al., 2024, Wu et al., 2024] to improve the performance.
4.3.2 Effect of number of synthetic data:
The results of the proposed method’s experiments with different amounts of generated synthetic data are summarized in Table 3, R/S refers to the number of real/synthetic images. The metrics we use to evaluate data imbalance include Entropy, and Class Imbalance Ratio (CIR). First, when not training on synthetic data, the result is 46.54 mIoU. When training the model using synthetic data with a quantity of approx $|D_{origin}|$, approx $2\times|D_{origin}|$, and approx $3\times|D_{origin}|$, the data balance metrics remain stable at a better level than before; the model performance is also better. We found that the amount of synthetic data approx $3\times|D_{origin}|$ has affected the model performance when the performance is now lower than when the number is approx $1\times|D_{origin}|$, approx $2\times|D_{origin}|$.
4.3.3 Text prompt seclection:
Table 4 compares different text prompt selection methods for generative modeling. We compare the performance of three prompt types: Generated caption generated from the image captioning model, Simple text prompt listing the classes in the image, and Class-Prompt Appending, a combination of the two prompt types. Class-Prompt Appending outperforms the other two methods by 49.72, precisely 3.20 and 1.80 better than Generated caption and Simple text prompt, respectively, in mIoU. These results show that the Class-Prompt Appending text prompt selection method can support SD in generating diverse datasets and ensuring accurate attention.
4.3.4 Effect of different trade-off scales:
@@
4.3.5 Other Visual Priors
We compare using different visual priors for both T2I-Adapter [Mou et al., 2023] and ControlNet [Zhang et al., 2023]: Line Art [Chan et al., 2022], Canny [Canny, 1986], Sketch [Su et al., 2021]. We also combine Inpainting [Rombach et al., 2022] with our methods (excluding Visual Prior Combination). Although the T2I-Adapter combined with Line Art gives the best result at 49.72% mIoU, other visual priors also show competitive performance, especially Canny on both T2I-Adapter and ControlNet. Figure 5 shows some imaging results with controllable diffusion model.
Figure 5: Some image generation results from various methods when combined with our proposed approach.
Table 6: Different visual priors controlled the enhancement results. All used the Stable Diffusion XL model.
@@ 4.3.5 Other Visual Priors
mIoU (%) 49.72 48.95 47.52 48.95 47.56
@@ 5.1 Limitations
While our approach effectively generates synthetic images to augment data for semantic segmentation, there are certain limitations to consider. First, the results in Tables 1, 3 show that the model’s performance may decrease when the number of synthetic images is large or more significant than the number of original real images. This may be because the synthetic images do not completely guarantee the original images’ location and quantity of labels. Additionally, as the images produced by the Stable Diffusion [Podell et al., 2024] model are trained on the LION-5B dataset [Schuhmann et al., 2022], the resulting images do not share the same distribution as the target dataset. So, the synthetic data cannot completely replace the original training dataset used to train the model.
@@ 5.2 Conclusion
We introduce a data augmentation pipeline based on a controllable diffusion model for semantic segmentation tasks. Our evaluations were performed on the PASCAL VOC dataset, and the results show that our pipeline significantly improves the performance of semantic segmentation models. We note that our method can be combined with other data augmentation methods to enhance performance further. This significant advancement opens the door for future research into data augmentation using generative models.

### v3

@@ 3.3 Create class-balancing dataset
In the default setting, we choose an appropriate $n_{balance}$ such that $|D_{gen}|\approx|D_{origin}|$ where |||| . |||| is the number of images in the dataset. In Section 4.3.2, we further discuss the impact of varying the number of generated synthetic images.
@@ 4.1.1 Dataset
In this section, we evaluate our method on the segmentation datasets VOC7 and VOC12 [Everingham et al., 2015]. PASCAL VOC 2007 has 422 images annotated for semantic segmentation, split into 209 training and 213 validation images. Meanwhile, VOC12 has training and validation sets, including 1.464 and 1.449 images. In addition to training on the entire VOC12 dataset (1.464 images), we also train our model using 1/2 (732 images), 1/4 (366 images), 1/8 (183 images), and 1/16 (92 images) partition protocols [Wang et al., 2022]. These evaluations on smaller subsets demonstrate the effectiveness of our method in real-world, limited-data scenarios.
@@ 4.1.2 Implementation details
We construct our framework on the deep learning PyTorch framework [Paszke et al., 2019] and T2I-Adapter [Mou et al., 2023] using Stable Diffusion XL 1.0 [Podell et al., 2024] with 30 time steps. We generate data using values for $\omega_{1}$ and $\omega_{2}$, as defined in Section 4.3.4. For semantic segmentation, we employ the DeepLabV3+ [Chen et al., 2018], PSPNet [Zhao et al., 2017], and Mask2Former [Cheng et al., 2022] with segmenters implemented in the MMSegmentation framework [MMSegmentation Contributors, 2020]. We utilize the SGD optimizer with standard settings in MMSegmentation. We train our models with an input image size of 512$\times$512 with 30k steps on the Pascal VOC datasets, including VOC7 and VOC12. During training, we only apply simple transformation methods to augment data, such as: RandomResize, RandomCrop, RandomFlip. The performance of the trained model on the applied data is assessed using the Mean Intersection over Union (mIoU) metric.
@@ 4.2.1 Quantitative results:
Table 2: Evaluation results of the DeepLabV3+ model on the PASCAL VOC7 dataset trained on $D_{origin}$ and $D_{final}$.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$D_{origin}$ 89.32 75.35 43.90 53.95 38.03 51.72 50.81 67.41 72.0 7.77 24.61
$D_{final}$ 88.69 67.67 45.35 57.00 38.07 56.08 70.14 70.67 73.46 28.81 45.45
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] mIoU (%)
$D_{origin}$ 41.67 45.16 42.29 66.08 72.71 42.00 14.07 24.68 40.74 12.96 46.54
$D_{final}$ 39.48 36.27 50.10 55.80 71.49 40.29 21.91 18.10 48.13 32.72 50.27
Figure 5: The mIoU (%) of the DeepLabV3+ model over 30k training steps. The star symbol indicates the point of convergence, where the model achieves its highest performance on the validation set.
The results presented in Table 1 demonstrate that combining augmented data ($D_{gen}$) with the original dataset ($D_{origin}$) improves the performance of various segmentation models on the VOC7 and VOC12 datasets. All models, including DeepLabV3+, PSPNet, and Mask2Former, significantly improve when augmented data. DeepLabV3+ consistently performs the best across different dataset sizes. We note that as the amount of real-world data increases, the accuracy of the generated images and labels becomes crucial. Mismatched generated images in the synthetic data can lead to performance degradation; this is observed in Mask2Former when trained on the VOC12 dataset with 1464 images.
Figure 6: Some results show the limitations of direct generation and the effectiveness of our method to overcome them.
Table 3: Impact of Class-Prompt Appending (1), Visual Prior Blending (2), Class balancing algorithm (3), and Post Filter [Fang et al., 2024] (4).
@@
$\checkmark$ $\checkmark$ $\checkmark$ 50.27
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 52.23
@@
69 209/845 4.057 0.225 47.32
Figure 5 visualizes the validation set accuracy on the PASCAL VOC7 dataset over 30,000 steps using the DeepLabV3+ model. The results show that training the model on the $D_{final}$ achieves 50.27% mIoU, 3.73% higher than training solely on the $D_{origin}$. The visualization also indicates that the model converges earlier when trained on the $D_{origin}$ compared to the $D_{final}$. Additionally, the detailed performance for each class in the PASCAL VOC dataset provided in Table 2 demonstrates that most classes show improved accuracy when trained on the $D_{final}$ dataset. Notably, some classes, such as “bus”, “cow”, “chair”, and “TV monitor”, exhibit significant improvements.
4.2.2 Qualitative results
In Figure 6, each row presents three images: the original image, the image generated by the Controllable Generation model (using prompts generated by BLIP-2), and the image produced by the generation model with our proposed. For the images generated by the SD model without our method, the first two rows illustrate cases where labeled classes in the image are missing in the generated description, causing the model to fail in accurately generating all those classes. In the third row, the image contains objects belonging to the “cow” class with smaller sizes, which prevents the Image Captioning Model from including “cow” in the description, resulting in an inaccurate generated image without cows. In the last row, although the case is relatively simple and the prompt includes all the labeled classes, the generated image still fails to depict the structure of both “bird” objects accurately. All four images are significantly improved with our proposed pipeline, with the labeled objects entirely generated and their structure well preserved. This demonstrates that our method effectively addresses the limitations of directly using Controllable Image Generation models.
4.3 Ablation Study
4.3.1 Effects of different methods:
The performance of the proposed methods is summarized in Table 3. When using the Baseline with generated prompts, the performance is only as low as 42.15% mIoU, which is lower than when training on the original data. This demonstrates the essential nature of the proposed methods when generating synthetic data. Our method yields a 50.27% mIoU, showing its effectiveness. Additionally, we experimented with combining the Post Filter with the Category-Calibrated CLIP Rank [Fang et al., 2024] for generating synthetic data. The result of this combination is higher, which shows that our method can combine filters from previous studies [Fang et al., 2024, Wu et al., 2024] to improve the performance.
4.3.2 Effect of number of synthetic data:
@@
50.27
@@
The results of the proposed method’s experiments with different amounts of generated synthetic data are summarized in Table 4, R/S refers to the number of real/synthetic images. The metrics we use to evaluate data imbalance include Entropy, and Class Imbalance Ratio (CIR). Initially, without using synthetic data, the model achieves an mIoU of 46.54. When training with synthetic data in quantities of approximately $|D_{origin}|$, $2\times|D_{origin}|$, and $3\times|D_{origin}|$, the data balance metrics stabilize at a better level, and model performance improves. However, we observed that using synthetic data around $3\times|D_{origin}|$ negatively impacts performance, resulting in a decrease compared to training with $1\times|D_{origin}|$ and $2\times|D_{origin}|$ of synthetic data.
4.3.3 Text prompt seclection:
Table 5 compares different text prompt selection methods for generative modeling. We compare the performance of three prompt types: Generated caption generated from the Image Captioning model, Simple text prompt listing the classes in the image, and Class-Prompt Appending, a combination of the two prompt types. Class-Prompt Appending outperforms the other two methods by 50.27 mIoU (%), precisely 3.03 and 2.06 better than generated caption and simple text prompt, respectively, in mIoU. These results show that the Class-Prompt Appending text prompt selection method can support SD in generating diverse datasets and ensuring accurate attention.
Figure 7: Some image generation results from various Controllable models when combined with our proposed approach.
Table 7: Different visual priors controlled the enhancement results. All used the Stable Diffusion XL version.
@@ 4.3.3 Text prompt seclection:
mIoU (%) 50.27 48.95 47.52 48.95 47.56
4.3.4 Effect of different trade-off scales:
4.3.5 Other Visual Priors
We compare using different visual priors for both T2I-Adapter [Mou et al., 2023] and ControlNet [Zhang et al., 2023]: Line Art [Chan et al., 2022], Canny [Canny, 1986], Sketch [Su et al., 2021]. We also combine Inpainting [Rombach et al., 2022] with our methods (excluding Visual Prior Blending). Although the T2I-Adapter combined with Line Art gives the best result at 50.27% mIoU, other visual priors also show competitive performance, especially Canny on both T2I-Adapter and ControlNet. Figure 7 shows some image results with Controllable Diffusion model.
@@ 5.1 Limitations
While our approach effectively generates synthetic images to augment data for semantic segmentation, there are certain limitations to consider. First, the results in Tables 1 and 4 show that the model’s performance may decrease when the number of synthetic images is large or more significant than the number of original real images. This may be because the synthetic images do not completely guarantee the original images’ location and quantity of labels. Additionally, as the images produced by the Stable Diffusion [Podell et al., 2024] model are trained on the LION-5B dataset [Schuhmann et al., 2022], the resulting images do not share the same distribution as the target dataset. So, the synthetic data cannot completely replace the original training dataset used to train the model.
@@ 5.2 Conclusion
In this study, we introduced a novel data augmentation pipeline for semantic segmentation tasks based on Controllable Diffusion models. Our proposed methods, including Class-Prompt Appending, Visual Prior Blending, and a class-balancing algorithm, effectively address challenges associated with generating synthetic images while preserving the structure and class balance of labeled datasets. By combining synthetic and real-world data, we demonstrated improvements in segmentation performance on the PASCAL VOC datasets in terms of mIoU, compared to training on original datasets alone. These results validate the effectiveness of our approach, particularly in scenarios with limited data availability. Furthermore, our method can be seamlessly combined with other augmentation methods to further enhance performance. Through extensive experiments, we have demonstrated the versatility and robustness of our approach, providing a strong foundation for future research in data augmentation.